If Jane were to increase her study time from three hours to four hours a night, her grade point average would increase from 2.5 to 3.2. Calculate Jane's elasticity of grade point average with respect to hours studying and interpret your resul
1
Expert's answer
2019-04-25T09:55:28-0400

Jane's elasticity of grade point average with respect to hours studying is:

E = (3.2 - 2.5)/(4 - 3)×(4 + 3)/(3.2 + 2.5) = 0.7/1×7/5.7 = 0.86.

So, the relationship is inelastic.